150g of a gas at 10.0°C and 1.0 atm pressure occupies 913L volume. The volume of the gas increases to 988L and its temperature increases to 41.5°C while the pressure remains constant. Calculate ΔE, q and w for the gas. (Specific heat of the gas = 5.2J/g°C and 1L.atm = 101.3J). [4 marks]
1
Expert's answer
2020-06-08T15:48:23-0400

ΔE = q + w - the first law of thermodynamics.

"q=c\cdot m\cdot\Delta T=5.2J/g°C\cdot150g \cdot(41.5°C-10°C)=24570J"

"w=-p\cdot \Delta V=-1atm\cdot(988L-913L)=-75L\cdot atm=-75\cdot 101.3J=-7597.5J\approx7598J"

"\Delta E=q+w=24570J-7598J=16972J"


Answer:

"\Delta E=16975J"

"q=24570J"

"w=7598J"